if each student in a class sends a card to every one of his class mate the total number of cards sent out will be rs 1560. find the number of student in the class.
Answer by checkley77(12844)

Let x be the number of students in the class
Let (x-1) be the number of cards each student sends (everyone but themselves)
x(x-1)=1,560
x^2-x-1,560=0
(x-40)(x+39)=0
x-40=0
x=40 students are in the class.
Proof:
40*39=1,560
1,560=1,560